OpenStudy (smlohr):

An 8.5g ice cube is placed into 255g of water. Calculate the temperature change in the water upon the complete melting of the ice. Assume the all of the energy required to melt the ice comes from the water.

OpenStudy (sahar009):

The heat fusion of ice is 333 J/g. Q = 8.5 * 333 = 2830.5 J This specific heat of water is 4.186 J/(g * C) For water, Q = 255 * 4.186 * = 1067.43 * ∆T Set this equal to 2830.5 and solve for * ∆T 1067.43 * ∆T = 2830.5 ∆T = 2830.5 ÷ 1067.43 This is approximately 2.65˚ C